# Beyond Static AI Evaluations: Advancing Human Interaction Evaluations for LLM Harms and Risks

### v2

@@ 1 Introduction
Artificial intelligence (AI) model evaluations have become central in developers and regulators’ efforts to ensure that AI models are “safe.” Governments like that of the United Kingdom, through its AI Safety Institute (UK Government 2024), and the United States, through a recent Executive Order, emphasize the importance of conducting model evaluations for various risks like discrimination and cybersecurity (The White House 2023); AI labs, including OpenAI with its Preparedness Framework and Anthropic with its Responsible Scaling Policy (OpenAI 2023; Anthropic 2023), propose utilizing model evaluations to monitor and mitigate misuse and catastrophic risks; and, academic researchers are developing safety evaluation datasets at unprecedented rates (Röttger et al. 2024). This has positioned model evaluations as integral to a range of important decisions on the safe development and deployment of AI systems.
@@
While drawing on the history of research in natural language processing (NLP), the evaluation of general purpose AI models faces new challenges. These models have open-ended input and output spaces, and thus can produce unpredictable and varied responses which make designing comprehensive evaluations challenging (Ganguli et al. 2022). Their model evaluation results are also sensitive to different prompting strategies (Liang et al. 2022), suffer from replicability issues (Ganguli et al. 2023; Hutson 2018), and can be compromised by data leakage from test sets (Jiang et al. 2024). These challenges have prompted calls for intensified and interdisciplinary research efforts to enhance the robustness, reliability, and replicability of evaluations (Chang et al. 2024).
@@
Current safety evaluation approaches primarily develop datasets and benchmarks to evaluate model completions for hazardous behaviors and capabilities (Sharma et al. 2023; Parrish et al. 2021; Wang et al. 2023). While such approaches are important in identifying potential downstream harms, they insufficiently address the ‘sociotechnical gap’ — the discrepancy between what safety assessments predict in controlled, model-only settings and how models actually perform in the environments in which they are deployed (Weidinger et al. 2023; Liao and Xiao 2023; Raji et al. 2022). By assessing models in isolation, these evaluation approaches incorporate human factors in limited ways, thereby failing to capture the complexity of real-world human-model interaction (Chang et al. 2024).
@@
This sociotechnical gap manifests across three dimensions: (1) joint performance: while many AI use cases involve human interaction, most benchmarks are non-interactive and might not reveal capabilities and behaviors that may only appear through human-model collaboration and possible feedback effects (Pan et al. 2024); (2) evaluation tasks: analyses of real-world usage data reveal significant discrepancies between benchmark tasks and those performed in practical scenarios, indicating that current evaluations may not accurately reflect model behaviors outside of ‘lab settings’ (Ouyang et al. 2023); and (3) human impact: current evaluations are insufficient for examining the direct impact of model capabilities and behaviors on users (Weidinger et al. 2023).
@@
Recognizing these limitations, there has been an increasing recognition of the need to involve human users in assessments of model risks. In this work, we provide evaluators — whether in labs, governments, or academia — with a framework to design and interpret human interaction evaluations (HIEs). We focus on large language models (LLMs) due to their ubiquity and influence in current applications, though we expect that many of our findings will also apply to other modalities (with additional considerations which we invite future work to address). Our key contributions are:
@@
  • Providing actionable recommendations. We conclude with recommendations for advancing the practice of HIEs (Section 6).
@@ 2.2 Reviewing the evaluation of human-LLM interaction
However, research on interactive evaluations of LLMs has been relatively limited, with some exceptions. Lee et al. notably present a novel approach extending LLM evaluations from static to interactive through three stages: (1) defining an interactive task, (2) constructing an interactive system, and (3) designing evaluation metrics (Lee et al. 2022). They evaluate several models across five tasks and find that better non-interactive performance does not always translate to better interactive performance, and that user judgment of performance can diverge from that of crowdsourced annotators. Other studies like Collins et al. utilize this framework to interactively evaluate three LLMs as assistants in proving undergraduate-level mathematics (Collins et al. 2023). User-led, interactive evaluations have also been explored as a way to compare the performance of different models, such as the popular use of Elo or relative win rates between models (Chiang et al. 2024).
Most existing work focuses on interaction as a source of feedback for improving model performance; our study expands on this work, notably Lee et al. (2022) and existing insights from HCI, to specifically improve model safety as opposed to general model performance (Subramonyam et al. 2024; Gordon et al. 2021).
@@ 3.1 Definition
Weidinger et al. (2023) first use the term “human interaction evaluation” to describe evaluations which study ”effects on people interacting with AI systems, and the human–AI dyad”. Here, we further refine and operationalize this definition. We define an HIE as the assessment of human-model interactions, which includes the process and/or the outcome of a human engaging with a model — that is, focusing on the mechanisms of interaction, the outcomes yielded, or a combination of the two.
Humans can be involved in model evaluations in different ways; in this paper, we distinguish between static evaluations and interactive evaluations. While static evaluations, like common dataset-based benchmarks, can still prompt models using human-written prompts and assess model outputs using annotations from human crowdworkers, interactive evaluations – which we focus on and develop here – engage humans as subjects of the evaluation who either respond to or actively elicit model outputs through interacting with the model (Chang et al. 2024; Lee et al. 2022).
@@ 3.2 Utility
We argue that HIEs can have greater internal and external validity compared to static evaluations, as well as increase the validity of static evaluations.
The internal validity of an evaluation is the extent to which the chosen datasets, survey questions, or prompts measure the target model properties in the controlled environment of the evaluation (Rooney et al. 2016; Raji et al. 2021). HIEs enable researchers to collect richer forms of observable data (Kiela et al. 2021). For example, researchers can observe the behaviors of the entire human-model system, including how participants’ beliefs change after interacting with the model, what they do with given information, and which model capabilities they leverage most frequently. Such detailed observations can be used to generate insights or create datasets, prompts, or other proxies that can more faithfully measure the target model properties (Mialon et al. 2023). Therefore, HIEs may not only be a possible substitute for some static evaluations, but may also be an important complement that tests the validity of static evaluations.
The external validity of an evaluation is the extent to which the insights gained from it apply beyond the specific controlled environment of the evaluation. HIEs allow for experimenting and drawing conclusions in contexts that are more akin to real-world contexts, as they directly involve human users interacting with models. This increases the likelihood that evaluation results generalize to real-world settings. For example, an HIE could assess the capacity of human users to elicit a particular model capability and identify the cases in which this occurs. HIEs are thus particularly useful when it comes to assessing not only the upper bound of capabilities (e.g., in the case of human uplift evaluations with domain experts) but also the extent to which everyday human-model interactions elicit model capabilities approaching this upper bound. Importantly, the representativeness of an HIE study sample is also a determiner of external validity as it concerns whether HIE results on a study population (e.g., a representative sample of U.S. adults crowdsourced from an online platform) can generalize to the target population (e.g., the population of the U.K. or the entire user base of a product) (Susukida et al. 2017).
@@
First, HIEs enhance the external validity of human impact assessment. Static evaluations often rely on grading a set of model outputs using simple metrics, assuming these are adequate proxies for human impact. However, this does not establish a direct connection to actual effects on human behavior and decision-making. For example, static benchmarks cannot directly assess the psychological impact of models on human users’ emotions, cognitions, and behaviors. This implicates harms like those from anthropomorphic language (Abercrombie et al. 2023), overreliance on models (Xu, Feng, and Chen 2023), model-driven deception and persuasion (Hackenburg et al. 2023; Karinshak et al. 2023a), and the engagement in friendships and romantic relationships with AI systems (Brandtzaeg, Skjuve, and Følstad 2022) — all of which HIEs may be more suited to address.
Second, many risks are specifically realized through human interaction, and are difficult to discover and study without additional human involvement. For example, static benchmarks do not fully capture effects of feedback loops of human-model influence, which previous work has shown can drive harmful behavior like in-context reward hacking (Pan et al. 2024). Conversely, the risks of a model which static benchmarks show can produce falsehoods, for example, may not be realized if humans consistently observe and reject the outputs as misleading.
@@
Furthermore, integrating the results of HIEs with additional contextual information can provide a clearer view of how, if at all, these risks would materialize at the societal level. For example, the feasibility of human interaction risks can be contingent upon the availability of resources such as time, money, and personnel. Consider the assessment of disinformation risks; it is crucial to consider both the findings of a relevant HIE and the effectiveness of existing social media defenses, like whether content is adequately watermarked or how recommendation systems respond to AI-generated content, to design assessments that more accurately reflect broader societal impact (Kapoor and Narayanan 2023).
@@ 3.3 Methods
HIEs can either be controlled, where interactions are set in a structured, lab-like setting to systematically study specific variables under predetermined conditions, or grounded, where interactions occur in naturalistic, real-world settings to observe how humans interact with models in real-life environments. We briefly recap the methods for conducting HIEs, organizing them around established methods in HCI (Kuniavsky 2003):
@@ 4.1 Patterns in emerging human interaction evaluations
A recent review by Weidinger et al. (2023) shows that the number of HIEs is limited but growing. We conduct a brief review of 21 such evaluations. We restrict our review to studies which utilize one of the methods from Section 3.3 to assess an aspect of model safety, risks, or harms. Our review consists of 11 studies from the repository of safety evaluations in Weidinger et al. (2023)’s review (last updated in December 2023), along with 10 additional studies which we add from existing knowledge and suggestions made by the paper authors. We examine these studies for common patterns and limitations, categorizing them by (1) risk/harm focus area, (2) methods, and (3) design attributes, namely target study population, nature of risks, interaction tasks, and number of models evaluated (as seen in Appendix Table 1). The insights from this review lay the groundwork for our framework, and are summarized below:
@@
Recent work has drawn attention to the importance of assessing marginal rather than absolute risks that stem from AI systems (Kapoor et al. 2024). We find that while earlier studies often focused on absolute risks — assessing the direct likelihood and severity of harms from model use, more recent studies tend to examine marginal risks, comparing the risks of models against a human baseline or that of other existing technologies.
@@
Research indicates that different user groups experience computing harms differently, highlighting the importance of evaluating diverse user interactions and perceptions when assessing model risks (Aizenberg and van den Hoven 2020). While around half of the studies examined effects across user groups, these examinations often occurred only after data collection. We find that few studies targeted a specific population, like intentionally using a representative sample, or otherwise justified their sampling strategy.
@@
The broader applicability of findings across different models and tasks is important for conducting comprehensive evaluations. Most studies assessed either a single model or system (e.g., a specific model within ChatGPT) or a single type of task (e.g., information retrieval), limiting their generalizability.
@@ 4.2 Framework
Synthesizing the results of our review with evaluation considerations from NLP and HCI, we propose a three-staged approach to constructing safety-focused HIEs: (1) identifying risk and/or harm area, (2) characterizing the use context, and (3) choosing the evaluation parameters.  [^3] While the design considerations outlined in our framework are useful for both controlled and grounded studies, grounded studies conducted in real-world environments (e.g., workplaces, schools) require additional environment considerations which are beyond the scope of this paper.
@@
Dicrimination risks Risks of model use leading to unfair treatment based on protected classes (e.g., race, age, gender) Comparison of discrimination risks from model use in decision-making vs. human decision-making Discrimination risks that remain after use of anti-discrimination measures (e.g., model fairness checks)
@@
What is the harmful use scenario? In HCI research, user goals or objectives have been shown to shape how they engage with systems and thus influence the outcomes of these engagements (Subramonyam et al. 2024). Thus, here, we group harmful use scenarios according to the user’s objectives in the interaction. In each scenario, we also consider the affected parties of any harm caused by use. We illustrate in Table 2 the four main scenarios which we believe cover the most salient failure modes of current concern (Mitchell 2024).
Scenario Misuse Unintended harm: personal impact Unintended harm: external impact
@@
Example(s) Influence operations, cybersecurity attacks Exposure to harmful stereotypes in model output Decision-maker trusts inaccurate model judgment hurting decision-subject
@@
  • System dimensions: models could be augmented by system messages and/or scaffolding 555“Scaffolding” is a term used to describe structured support for AI models that improve their ability to perform certain tasks or expand their affordances and spheres of influence. with prompt engineering, planning and reasoning frameworks, or tools.  [^6] Given the current state of the field, system dimensions are fairly limited compared to model dimensions. Hence, in this paper, we mostly focus on model and user dimensions but in the future the scope of system dimensions will likely grow.
@@
  • Exposure: human observes or is exposed to model output (e.g., human reads a model-generated message).
@@
Figure 1: Taxonomy of human-LLM interaction modes. “Instruction” refers to the The figure illustrates different human-LLM interaction paths from an initial set of instructions to completing goal-oriented or open-ended tasks.
@@
Discussion QA, explanation, brainstorming
@@
Finally, after selecting the risk area to evaluate and identifying the set up, the third stage is choosing an evaluation target within that set up and selecting evaluation metrics to measure the chosen target.
What is the evaluation target? HIEs shift the evaluation target from model-only output to joint human-model output. HIEs also introduce a new target: the interaction trace. Evaluating an interaction trace is evaluating the process of human-model interaction towards completing a task; this may include assessing the contextual appropriateness of model responses, the adaptability of the model to human inputs, and the efficiency of achieving the intended outcome. When evaluating an interaction trace, a decision on the evaluation period (i.e., for how long to evaluate) must also be made as it can affect the evaluation results. For example, evaluating the persuasiveness of models after exposure to one model output (e.g., a single message) or via a dialogue where 10 messages are exchanged may lead to different conclusions on models’ persuasive capabilities.
@@
Interaction trace - user experience Subjective Metrics that survey users for responses relating to their interaction experience Usability metrics, user satisfaction surveys, psychometrically validated measures (e.g., Social Responsiveness Scale (SRS), Decision Regret Scale)

### v4

@@ 1 Introduction
Artificial intelligence (AI) model evaluations have become central in developers’ and regulators’ efforts to ensure that AI systems are “safe.” Governments like that of the United Kingdom, through its AI Safety Institute (UK Government 2024), and the United States, through a recent Executive Order, emphasize the importance of conducting model evaluations for various risks like discrimination and cybersecurity (The White House 2023); AI labs, including OpenAI with its Preparedness Framework and Anthropic with its Responsible Scaling Policy (OpenAI 2023; Anthropic 2023), propose utilizing model evaluations to monitor and mitigate misuse and catastrophic risks; and, academic researchers are developing safety evaluation datasets at unprecedented rates (Röttger et al. 2024). This has positioned model evaluations as integral to a range of important decisions on the safe development and deployment of AI systems.
@@
While drawing on the history of research in natural language processing (NLP), the evaluation of current general purpose AI models faces new challenges. These models have open-ended input and output spaces, and thus can produce unpredictable and varied responses which make designing comprehensive evaluations challenging (Ganguli et al. 2022). Their model evaluation results are also sensitive to different prompting strategies (Liang et al. 2022), suffer from replicability issues (Ganguli et al. 2023; Hutson 2018), and can be compromised by data leakage from test sets (Jiang et al. 2024). These challenges have prompted calls for intensified and interdisciplinary research efforts to enhance the robustness, reliability, and replicability of evaluations (Chang et al. 2024).
@@
Current safety evaluation approaches primarily develop datasets and benchmarks to evaluate model completions for hazardous behaviors and capabilities (Sharma et al. 2023; Parrish et al. 2021; Wang et al. 2023). While such approaches are important in identifying potential downstream harms, they insufficiently address the ‘sociotechnical gap’ — the discrepancy between what safety assessments predict in controlled, model-only settings and how models actually perform in the environments in which they are deployed (Weidinger et al. 2023; Liao and Xiao 2023; Raji et al. 2022). By assessing models largely in isolation, these evaluation approaches incorporate human factors in limited ways, thereby failing to capture the complexity of human-model interactions (Chang et al. 2024).
@@
This sociotechnical gap manifests across three dimensions: (1) joint performance: while many AI use cases involve human interaction, most benchmarks are non-interactive and might not reveal capabilities and behaviors that may only appear through human-model collaboration and possible feedback effects (Pan et al. 2024); (2) evaluation tasks: analyses of real-world usage data reveal significant discrepancies between benchmark tasks and those performed in practical scenarios, indicating that current evaluations may not accurately reflect model behaviors outside of ‘lab settings’ (Ouyang et al. 2023); and (3) human impact: current evaluation approaches are insufficient for operationalizing complex constructs and examining the direct impact of model behaviors on human users (Weidinger et al. 2023).
@@
Recognizing these limitations, there has been an increasing recognition of the need to expand how safety evaluations incorporate human factors. In this work, we motivate the need for this expansion and provide model evaluators — whether in labs, governments, or academia — with a framework to design human interaction evaluations (HIEs). We focus on large language models (LLMs) due to their ubiquity and influence in current applications, though we expect that many of our findings will also apply to other modalities (with additional considerations which we invite future work to address). Our key contributions are:
@@
  • Providing actionable recommendations. We conclude with recommendations for advancing HIEs (Section 6).
@@ 2.2 Reviewing the evaluation of human-LLM interaction
However, research on evaluations of human-LLM interactions has been relatively limited, with some exceptions. Lee et al. notably present a novel approach extending LLM evaluations from static to interactive through three stages: (1) defining an interactive task, (2) constructing an interactive system, and (3) designing evaluation metrics (Lee et al. 2022). They evaluate several models across five tasks and find that better non-interactive performance does not always translate to better interactive performance, and that user judgment of performance can diverge from that of crowdsourced annotators. Other studies like Collins et al. utilize this framework to interactively evaluate three LLMs as assistants in proving undergraduate-level mathematics (Collins et al. 2023). User-led, interactive evaluations have also been explored as a way to compare the performance of different models, such as the popular use of Elo or relative win rates between models (Chiang et al. 2024).
Most existing work focuses on interaction as a source of feedback for improving model performance; our study expands on this work, notably Lee et al. (2022) and existing insights from HCI, to specifically understand and improve model safety as opposed to general model performance (Subramonyam et al. 2024; Gordon et al. 2021).
@@ 3.1 Definition
Weidinger et al. (2023) first use the term “human interaction evaluation” to describe evaluations which study ”effects on people interacting with AI systems, and the human–AI dyad.” Here, we further refine and operationalize this definition. We define an HIE as the assessment of human-model interactions, which includes the process and/or the outcome of a human engaging with a model — that is, focusing on the mechanisms of interaction, the outcomes yielded, or a combination of the two.
Humans can be involved in building model evaluations in different ways; in this paper, we distinguish between static evaluations and interactive evaluations. While static evaluations, like common dataset-based benchmarks, can still prompt models using human-written prompts and assess model outputs using annotations from human crowdworkers, interactive evaluations – which we focus on and develop here – engage humans as subjects of the evaluation who either respond to or actively elicit model outputs through interacting with the model (Chang et al. 2024; Lee et al. 2022).
@@ 3.2 Utility
We argue that for some target constructs, HIEs can have greater internal and external validity compared to static evaluations, and that they may also increase the validity of static evaluations.
The internal validity of an evaluation is the extent to which the chosen datasets, survey questions, or prompts measure the target model properties in the controlled environment of the evaluation (Rooney et al. 2016; Raji et al. 2021). HIEs enable researchers to collect richer forms of observable data (Kiela et al. 2021). For example, researchers can observe the behaviors of the entire human-model system, including how participants’ beliefs change after interacting with models, what they do with given information, and which model capabilities they leverage most frequently. Such detailed observations can be used to generate insights or create datasets, prompts, or other proxies that can more faithfully measure the target model properties (Mialon et al. 2023). Therefore, HIEs may not only be a possible substitute for some static evaluations, but may also be an important complement that validates static evaluations.
The external validity of an evaluation is the extent to which the insights gained from it apply beyond the specific controlled environment it is conducted in. HIEs allow for experimenting and drawing conclusions in contexts that are more akin to real-world, naturalistic contexts, as they directly involve human users interacting with models. This can increase the likelihood that evaluation results generalize to real-world settings. For example, an HIE could assess the capacity of human users to elicit a particular model capability and identify the cases in which this occurs. HIEs are thus particularly useful when it comes to assessing not only the upper bound of capabilities (e.g., in the case of red-teaming with domain experts) but also the extent to which everyday human-model interactions elicit model capabilities approaching this upper bound. Importantly, the representativeness of an HIE study sample is also a determiner of external validity as it concerns whether HIE results on a study population (e.g., a representative sample of U.S. adults crowdsourced from an online platform) can generalize to the target population (e.g., the population of the U.K. or the entire user base of a product) (Susukida et al. 2017).
@@
First, HIEs enhance the external validity of human impact assessment. Static evaluations typically grade model outputs using specific metrics, assuming these metrics adequately represent human impact, without directly measuring the actual impact on users. For example, static evaluations cannot directly assess the psychological impact of models on human users’ emotions, cognitions, and behaviors. This implicates harms like those from anthropomorphic language (Abercrombie et al. 2023), overreliance on models (Xu, Feng, and Chen 2023), model-driven deception and persuasion (hackenburgcomparing; Karinshak et al. 2023a), and the engagement in friendships and romantic relationships with AI systems (Brandtzaeg, Skjuve, and Følstad 2022) — all of which HIEs may be more suited to address as a starting point.
Second, many risks are specifically realized through, often multi-turn, human interaction, and are difficult to discover and study without additional human involvement. For example, static evaluations do not fully capture effects of feedback loops of human-model influence, which previous work has shown can drive harmful behavior like in-context reward hacking (Pan et al. 2024). Conversely, the risks of a model which static evaluations show can produce falsehoods, for example, may not be realized if humans consistently observe and reject the outputs as misleading.
@@
Furthermore, integrating the results of HIEs with additional contextual information can provide a clearer view of how, if at all, these risks would materialize at the societal level. For example, the feasibility of human interaction risks can be contingent upon the availability of resources such as time, money, and personnel. Consider the assessment of disinformation risks; it is crucial to incorporate both the findings of a relevant HIE and the effectiveness of existing social media defenses, like whether content is adequately watermarked or how recommendation systems respond to AI-generated content, to design assessments that more accurately reflect broader societal impact (Kapoor and Narayanan 2023).
@@ 3.3 Methods
HIEs can either be controlled, where interactions are set in a structured, lab-like setting to systematically study specific variables under predetermined conditions, or grounded, where interactions occur in naturalistic, real-world settings to observe how humans interact with models in real-life environments. We briefly recap the methods for conducting HIEs, organizing them around established methods in HCI and adjacent fields (Kuniavsky 2003). Each method has different strengths in extracting various insights on human-LLM interaction:
@@ 4.1 Patterns in emerging human interaction evaluations
A recent review by Weidinger et al. (2023) shows that the number of HIEs is limited but growing. We conduct a brief review of 21 such evaluations. We restrict our review to studies which utilize one of the methods from Section 3.3 to assess an aspect of model safety, risks, or harms. Our review consists of 11 studies from the repository of safety evaluations in Weidinger et al. (2023)’s review (last updated in December 2023), along with 10 additional studies which we add from existing knowledge and suggestions made by the paper authors. We examine these studies for common patterns and limitations, categorizing them by (1) risk/harm focus area, (2) methods, and (3) design attributes, namely target study population, nature of risks, interaction tasks, and number of models evaluated (details in Appendix Table 1). The insights from this review lay the groundwork for our framework, and are summarized below:
@@
Recent work has drawn attention to the importance of assessing marginal rather than absolute risks from AI system use (Kapoor et al. 2024). We find that while earlier studies often focused on absolute risks — assessing the direct likelihood and severity of harms from model use, more recent studies tend to examine marginal risks, comparing the risks of models against a human baseline or that of other existing technologies.
@@
Research indicates that different user groups experience computing harms differently, highlighting the importance of evaluating diverse user interactions and perceptions when assessing model risks (Aizenberg and van den Hoven 2020). While around half of the studies examined effects across user groups, these examinations often occurred only after data collection. We find that few studies targeted a specific population, like intentionally using a representative sample, set of user personas, or otherwise justified their sampling strategy.
@@
The broader applicability of findings across different models and tasks is important for conducting relatively comprehensive evaluations. Most studies assessed either a single model or system (e.g., a specific model within ChatGPT) or a single type of task (e.g., information retrieval), limiting their generalizability and coverage.
@@ 4.2 Framework
Synthesizing the results of our review with evaluation considerations from NLP and HCI, we propose a three-staged approach to constructing safety-focused HIEs using human subject experiments: (1) identifying risk and/or harm area, (2) characterizing the use context, and (3) choosing the evaluation parameters.  [^3] While the design considerations outlined in our framework are useful for both controlled and grounded studies, grounded studies conducted in real-world environments (e.g., workplaces, schools) require additional environment considerations which are beyond the scope of this paper.
@@
Discrimination risks Risks of model use leading to unfair treatment based on protected classes (e.g., race, age, gender) Comparison of discrimination risks from model decision-making vs. human decision-making Discrimination risks that remain after use of anti-discrimination measures (e.g., model fairness checks)
@@
What is the harmful use scenario? In HCI research, user goals or objectives have been shown to shape how they engage with systems and thus influence the outcomes of these engagements (Subramonyam et al. 2024). Thus, here, we group harmful use scenarios according to the user’s objectives in the interaction. In each scenario, we also consider the affected parties of any harm caused by use. We propose in Table 2 four scenarios which we believe address some of the most salient failure modes of current concern (Mitchell 2024). Additionally, evaluations may be grounded in a specific use domain (e.g., medicine, education) to target domain-specific considerations.
Scenario Misuse / adversarial testing Unintended harm: personal impact Unintended harm: external impact
@@
Example(s) Influence operations, cybersecurity attacks, hate speech Exposure to harmful stereotypes in model output Decision-maker trusts inaccurate model judgment hurting decision-subject
@@
  • System dimensions: models could be augmented by system messages and/or scaffolding 555“Scaffolding” is a term used to describe structured support for AI models that improve their ability to perform certain tasks or expand their affordances and spheres of influence. with prompt engineering, planning and reasoning frameworks, or tools. [^6] Given the current state of the field, system dimensions are fairly limited compared to model dimensions. Hence, in this paper, we mostly focus on model and user dimensions but in the future we expect the scope of system dimensions to grow.
@@
  • Exposure: human observes or is exposed to a single or discrete set of pre-produced model output (e.g., human reads a model-generated message).
@@
Figure 1: Taxonomy of human-LLM interaction modes. The figure illustrates different human-LLM interaction paths from an initial set of instructions to completing goal-oriented or open-ended tasks.
@@
Discussion explanation, brainstorming
@@
Finally, after selecting the risk area to evaluate and identifying the use context, the third stage is choosing an evaluation target within that set up and selecting evaluation metrics to measure the chosen target.
What is the evaluation target? HIEs shift the evaluation target from model-only output to joint human-model output. HIEs also introduce a new target: the interaction trace. Evaluating an interaction trace is evaluating the process of human-model interaction towards completing a task; this may include assessing the contextual appropriateness of model responses, the adaptability of the model to human inputs, and the efficiency of achieving the intended outcome over multiple turns of interaction. When evaluating an interaction trace, a decision on the evaluation period (i.e., for how long or how many turns to evaluate) must also be made as it can affect the evaluation results. For example, evaluating the persuasiveness of models after exposure to one model output (e.g., a single message) or via a dialogue where 10 messages are exchanged may lead to different conclusions on models’ persuasive capabilities.
@@
Interaction trace - user experience Subjective Metrics that survey users for responses relating to their interaction experience Usability metrics, user satisfaction surveys, psychometrically validated surveys and behavioral measures (e.g., Decision Regret Scale)